OpenStudy (anonymous):

Simplify (9y^7 / 15x^5) / (45y^3 / 5x^2) Simplify (9y^7 / 15x^5) / (45y^3 / 5x^2) @Mathematics

OpenStudy (asnaseer):

\[\frac{9y^{7}}{15x^{5}}/\frac{45y^{3}}{5x^{2}}=\frac{9y^{7}}{15x^{5}}*\frac{5x^{2}}{45y^{3}}=\frac{y^{4}}{15x^{3}}\]

OpenStudy (anonymous):

The answer I got was 27y^10 / 5x^7

OpenStudy (lalaly):

asnaseer got it right:)

OpenStudy (asnaseer):

remember that you can turn a fraction division into a fraction multiplication by inverting the divisor. so:\[\frac{a}{b}/\frac{c}{d}=\frac{a}{b}*\frac{d}{c}\]

OpenStudy (anonymous):

Oh my bad I meant this equation: (9y^7 / 15x^5) / ( 5x^2/45y^3 )

OpenStudy (asnaseer):

\[\frac{9y^{7}}{15x^{5}}/\frac{5x^{2}}{45y^{3}}=\frac{9y^{7}}{15x^{5}}*\frac{45y^{3}}{5x^{2}}=\frac{27y^{10}}{5x^{7}}\]

OpenStudy (anonymous):

Oh thank you So i did get it right? and Im soo sorry making you do that extra work before :/

OpenStudy (asnaseer):

no problem.